OpenStudy (anonymous):

4b-5=-t, solve for b

OpenStudy (lukecrayonz):

4b-5=-t +5 +5 4b=5-t /4 /4 b=5-t/4

OpenStudy (lukecrayonz):



OpenStudy (anonymous):